Moment of inertia of a semicircular ring of mass πkg and radius 2cm about the axis passing through a point perpendicular to it's plane
1
Expert's answer
2019-07-29T09:58:18-0400

Here the location of point is not given.

Assume the point is located at centre of mass which is located at position "d=\frac{2R}{\varPi}" from horizontal axis

Then MI is given as "MI=MR^2" =12.56 Kg-cm2

If the point is located at a point on horizontal axis,

Then MI is given as "MI=MR^2+Md^2"

Where,"d=2R/\varPi"

Thus

MI=3.14 x4+4/3.14

MI=12.56+1.27

MI=13.83 kg-cm2